A social club charters a plane at a cost of $34848 to take a group of members on an excursion to Atlantic City. At the last minute, 11 people in the group decide not to go. This raises the transportation cost per person by $24 . How many people originally intended to take the trip?
Answer by stanbon(75887)

A social club charters a plane at a cost of $34848 to take a group of members on an excursion to Atlantic City. At the last minute, 11 people in the group decide not to go. This raises the transportation cost per person by $24 . How many people originally intended to take the trip?
------------------------
Let "x" be the original number of trippers.
Average cost = 34848/x
----
Changed number of trippers: x-11
New average cost = 34848/(x-11)
-------------------------
Equation:
New Av - Old Av = 24
---
34848/(x-11) - (34848/x) = 24
11*34848 = 24x(x-11)
15972 = x^2-11x
---
x^2-11x-15972 = 0
(x+121)(x-132) = 0
---
Positive solution: x = 132 (original # of trippers)
======================================================
Cheers,
Stan H.